OpenStudy (anonymous):

Find the sum of the arithmetic sequence. -10, -7, -4, -1, 2, 5, 8 11 -70 0 -7

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Goshhh cant you add them just add them you get -7

OpenStudy (anonymous):

oh so all i do is add?

OpenStudy (anonymous):

that's what I'd do since they're listed

OpenStudy (anonymous):

what if they give me only 4 terms and tell me to find the sum of the first 9

OpenStudy (anonymous):

nevermind all you do is find the difference then find the first 9 terms then add them right?

OpenStudy (anonymous):

Find the sum of the first 9 terms of the sequence. Show all work for full credit. 2, -7, -16, -25, ...

OpenStudy (anonymous):

for that one they probably want to see you use the formulas. Can you come up with a equation for a_n using the formula?

OpenStudy (anonymous):

n-9?

OpenStudy (anonymous):

Well ye but if you should use equations same thing so d=+3 and they are 7 terms so n=7 S(7)=7/2 (-10+8)=7/2(2) =7

OpenStudy (anonymous):

this one. where you have to find d and a_1 \[a_n=a_1+d(n-1)\]

OpenStudy (anonymous):

Sorry -7*

OpenStudy (anonymous):

so an=2 + -9(n-1)?

OpenStudy (anonymous):

This is a straight arithmetic sequence guys no need for this d is simply d=Un+1 -Un

OpenStudy (anonymous):

yes. and they want the sum of the first 9 so now find a_9

OpenStudy (anonymous):

a9 would = -70 right?

OpenStudy (anonymous):

yes. Now for the sum use this formula \[S_n=\frac{ n }{ 2 }(a_1+a_n)\] so \[S_9=\frac{ n }{ 2 }(a_1+a_9)\]

OpenStudy (anonymous):

so the sum would =-306

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The sum of the first 9 is 18 S9=n/2 ( 2a1+(n-1)d = 9/2 (-20+24) =36/2 =18

OpenStudy (anonymous):

Oh sorry used another sequence

OpenStudy (anonymous):

@joyraheb different set of numbers from the sequence at the very top

OpenStudy (anonymous):

ok good because i was about to say, i thought i finally understood this then that messed me up hahaha

OpenStudy (anonymous):

but thank you!

OpenStudy (anonymous):

Oh ok so d=-9 so S9=-4.5(-14+90)= -342 right?

OpenStudy (anonymous):

no s9= (9/2)(2+(-70)) = -306

OpenStudy (anonymous):

because of this equation Sn=n/2(a1+an) @joyraheb